Calculate the capacitance of a capacitor whose plates are 25cm x 4.0 cm and are separated by a 2.0mm air gap.

1
Expert's answer
2022-02-02T05:15:19-0500

The formula to calculate the capacitance with dielectric is


"C=k\epsilon_0 \frac{A}{d}
\\ \text{}
\\ \text{where k is the dielectric constant, }\epsilon_0
\\ \text{is the absolute permittivity, A is the}
\\ \text{area of the plates, and d}
\\ \text{is the distance between them.}"


Since k = 1.00059 for air and "\epsilon_0"= 8.854 X 10-12 F/m, we proceed to substitute and we find:


"C = (1.00059)(8.854\times 10^{-12}\frac{F}{\cancel{m}})\frac{(0.25\cancel{\,m})(40.0 \cancel{\,mm})}{2.0 \cancel{\,mm}}
\\ \text{}
\\ \therefore C = 4.43\times 10^{-11}\,F"


In conclusion, the capacitance for this array is 4.43 X 10-11 F.